# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ 6.1 Synthetic Test Problems
Figure 3: Cumulative time cost of algorithms (in seconds).
Figure 4: Sensitivity analysis of hyperparameter $Z$ on GSSBO.
Figure 5: Sample distribution: no selection (left), gradient-based selection (middle), and random selection (right).
To assess the performance of our proposed methods, we test five benchmark functions, Eggholder2, Hart6, Levy20, Powell50, and Rastrigin100.
@@
Computational Efficiency Analysis. Figure 3 compares the cumulative runtime over $400$ iterations on low-dimensional Hart6 and high-dimensional Powell50 (results for other test functions are provided in the appendix due to space constraints). In both plots, VecchiaBO, SVI-GP, and LR-Firstm incur a rapidly accelerating runtime, whereas GSSBO and RSSBO remain notably lower than GP-UCB. While VecchiaBO reduces the cost of GP fitting by conditioning on nearest neighbors, its runtime is dominated by the costly maintenance of a structured neighbor graph, which scales poorly with dimensionality and sample size. SVI-GP introduces significant overhead from iterative optimization of pseudo-points and variational parameters via gradient descent, particularly in high dimensions. LR-Firstm, despite a fixed subset size $m$, incurs substantial costs from repeated maximization ordering and matrix factorizations as $n$ grows. These methods aim to reduce the computational complexity of GPs in the context of large budgets, but they overlook the substantial time costs associated with their approximation processes. In contrast, GSSBO and RSSBO cost much less time. Because we use a pilot-based method to compute gradients with minimal computational overhead, the fitting cost per iteration of GSSBO remains around $\mathcal{O}(M^{3})$ once we restrict the active subset to size $M\ll n$. The GSSBO and the random one are often similar in runtime, though the GSSBO can be slightly higher due to the overhead of the gradient-based sample selection process. By iteration 400, the GSSBO cuts the total runtime by roughly half compared to the Standard GP-UCB on both functions. This advantage of GSSBO increases more pronounced as $n$ increases.
@@
Optimization Performance Analysis. Figure 2 compares methods on multiple functions, evaluating cumulative regret. Overall, GSSBO achieves comparable performance with the Standard GP-UCB while outperforming the RSSBO. In low-dimensional problems such as Eggholder2 and Hart6, GSSBO has a subtle gap with Standard GP-UCB and VecchiaBO. In contrast, GSSBO significantly outperforms SVI-GP, LR-Firstm, and RSSBO. In high-dimensional settings, such as on Levy20, Powell50, and Rastrigin100, GSSBO achieves the smallest cumulative regret, surpassing VecchiaBO, SVI-GP and LR-Firstm. From the experimental results, we can observe that the cumulative of our algorithm is sublinear, which is consistent with the theoretical results. In particular, GSSBO achieves these results with a significant reduction in computation time, as shown in Figure 3. GSSBO strikes a combination of performance and efficiency in scalable optimization tasks by maintaining near-baseline regret while significantly improving computational efficiency.
@@
Sensitivity analysis of Hyperparameter $Z$. We further examine how the dynamic buffer parameter $Z$ affects our GSSBO. Figure 4 presents results on two functions: Hart6 and Powell50. For RSSBO, $Z$ remains fixed at 4, whereas for GSSBO, we vary $Z\in\{2,4,6,8\}$. On Hart6, larger $Z$ consistently boosts the GSSBO’s performance toward that of Standard GP-UCB, while the RSSBO lags in cumulative regret. Intuitively, for low dimensional problems, allowing the model to retain more samples helps preserve important information, bridging the gap with the Standard GP-UCB baseline. In contrast, in Powell50, smaller $Z$ leads to slightly better performances for GSSBO, reflecting the benefit of subset updates in high-dimensional landscapes. Sample selection maintains a robust exploration-exploitation balance. In general, low-dimensional tasks benefit from a larger $Z$, while high-dimensional problems perform better with more aggressive subset limiting, $Z$ can be effectively tuned to match the complexity of tasks.
@@ 6.2 Real-World Application
To assess the applicability of GSSBO in real-world applications, we applied Neural Architecture Search (NAS) to a diabetes-detection problem and used the diabetes dataset from the UCI repository (Dua & Graff, 2017). We modeled the problem of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of (batch size, Learning Rate, Learning Rate decay, hidden dim) in $[0,1]^{4}$ mapped to real hyperparameter ranges, where $y_{t}$ is the resulting classification error on the test set. Each iteration is repeated 5 times. The results in Figure 2 indicate that GSSBO outperforms all competitors.
6.3 Subset Samples Distribution Study
Figure 5 illustrates the sample distribution of GSSBO and Standard GP-UCB, on the first two dimensions of the Hartmann6 function. In this experiment, we recorded the first 200 sequential samples from a standard BO process and constrained the buffer size to 100. The objective is to identify the global minimum, and darker-colored samples correspond to values closer to the optimal solution.
During the optimization process, only a small number of samples are located near the optimal value (represented as the darker-colored samples being sparse). As shown in the middle panel, the gradient-based sample selection method selects a more informative and diverse subset, retaining more samples closer to the optimal or suboptimal, which is indicated by preserving a higher number of darker-colored samples in the figure. In contrast, the random selection strategy reduces the sample density uniformly across all regions, leading to a significant loss of samples near the optimal or suboptimal, as represented by the retention of many lighter-colored samples in the right panel. With gradient-based sample selection, the relative density of samples near the optimal and suboptimal regions increases, maintaining a more balanced distribution. This subset encourages subsequent iterations to focus on regions outside the optimal and suboptimal regions, promoting the exploration of other regions of the search space. As a result, the over-exploitation issue is mitigated.
@@ 7 Conclusion
Appendix A Theoretical Analysis
@@ A.1 Theorem 1:Analysis on subset GP:
Consider a GP model, where we assume $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime})),$ and given samples $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{t},y_{t})\}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, we have a noise model: $\mathbf{y}=f(\mathbf{X})+\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\sigma_{n}^{2}\mathbf{I}),$ where $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{t}]^{\top}\in\mathbb{R}^{t\times d}$ and $\mathbf{y}=[y_{1},y_{2},\ldots,y_{t}]^{\top}\in\mathbb{R}^{t}$. The posterior predictive distribution for a test point $\mathbf{x}_{*}$ is Gaussian with the following mean and variance: $\mu(\mathbf{x}_{*})=\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{y},$ $\sigma^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*},$ where $\mathbf{K}$ is the $t\times t$ kernel matrix evaluated at the training samples, i.e., $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$; $\mathbf{k}_{*}\in\mathbb{R}^{t}$ is the vector of covariances between the test point $\mathbf{x}_{*}$ and all training points, i.e., $(\mathbf{k}_{*})_{i}=k(\mathbf{x}_{*},\mathbf{x}_{i})$; $k_{**}=k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the kernel evaluated at the test point itself.
Sparse Approximation Using a Subset of Samples
@@
Instead of using all $N$ training samples, consider a subset samples $\mathcal{U}=\{\mathbf{u}_{1},\dots,\mathbf{u}_{M}\}$, where $M\ll N$. Define: $\mathbf{K}_{\mathcal{U}\mathcal{U}}\in\mathbb{R}^{M\times M},\quad\mathbf{K}_{\mathcal{D}\mathcal{U}}\in\mathbb{R}^{N\times M},\quad\mathbf{K}_{\mathcal{U}\mathcal{D}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}^{\top},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ is the kernel matrix among the $M$ inducing points, and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ represents the covariances between the full training samples in $\mathcal{D}$ and the inducing points in $\mathcal{U}$. Using Subset of Regressors (SoR), a low-rank approximation to the kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ is given by: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$ We can get the corresponding approximate posterior predictive mean using matrix decomposition and Woodbury formula for transformation.
Replacing $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ with $\hat{\mathbf{K}}$, the posterior predictive mean becomes: $\mu(\mathbf{x}_{*})=\mathbf{k}_{*}^{\top}\left(\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}\mathbf{y}.$ The Woodbury matrix formula states: $(\mathbf{U}\mathbf{V}+\mathbf{C})^{-1}=\mathbf{C}^{-1}-\mathbf{C}^{-1}\mathbf{U}(\mathbf{V}^{-1}+\mathbf{U}^{\top}\mathbf{C}^{-1}\mathbf{U})^{-1}\mathbf{U}^{\top}\mathbf{C}^{-1}.$ In our case: $\mathbf{C}=\sigma_{n}^{2}\mathbf{I}\in\mathbb{R}^{N\times N}$, $\mathbf{U}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\in\mathbb{R}^{N\times M}$, $\mathbf{V}=\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\in\mathbb{R}^{M\times M}$. We have: $\left(\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}=\sigma_{n}^{-2}\mathbf{I}-\sigma_{n}^{-4}\mathbf{K}_{\mathcal{D}\mathcal{U}}\left(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$ Substituting this result back into the posterior predictive mean:
$\tilde{\mu}(\mathbf{x}_{*})=\mathbf{k}_{*}^{\top}\left[\sigma_{n}^{-2}\mathbf{I}-\sigma_{n}^{-4}\mathbf{K}_{\mathcal{D}\mathcal{U}}\left(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}\right]\mathbf{y}$
$=\sigma_{n}^{-2}\mathbf{k}_{*}^{\top}\mathbf{y}-\mathbf{k}_{*}^{\top}\sigma_{n}^{-4}\mathbf{K}_{\mathcal{D}\mathcal{U}}\left(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}\mathbf{y}$
$=\sigma_{n}^{-2}\mathbf{k}_{*}^{\top}\mathbf{y}-\sigma_{n}^{-4}\mathbf{k}_{*\mathcal{U}}^{\top}\left(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I}\right)^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}\mathbf{y}$
$=\mathbf{k}_{*\mathcal{U}}^{\top}(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}\mathbf{y}.$
where $\mathbf{k}_{*\mathcal{U}}=\mathbf{k}_{*}^{\top}\mathbf{K}_{\mathcal{D}\mathcal{U}}$ represents the covariance vector between the test point $\mathbf{x}_{*}$ and the $M$ inducing samples.
In addition to the predictive mean, the exact posterior predictive variance of a GP is given by: $\sigma^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*}^{\top}(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*},$ where $k_{**}=k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the kernel evaluated at the test point. Under the sparse approximation (SoR), the approximate posterior predictive variance is:
$\tilde{\sigma}^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*\mathcal{U}}^{\top}(\mathbf{K}_{\mathcal{U}\mathcal{U}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{\mathcal{U}*},$
where $\mathbf{k}_{*\mathcal{U}}$ is the vector of covariances between the test point $\mathbf{x}_{*}$ and the $M$ inducing points, and $\mathbf{k}_{\mathcal{U}*}=\mathbf{k}_{*\mathcal{U}}^{\top}$.
@@
$\Delta\mu(\mathbf{x}_{*})=\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{y}.$
Define the following positive definite matrices: $\mathbf{A}=\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I},\quad\mathbf{B}=\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}.$ Then, the difference between $\mathbf{A}$ and $\mathbf{B}$ is: $\mathbf{A}-\mathbf{B}=\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}.$ We have $\|A^{-1}-B^{-1}\|\;\;\leq\;\;\|A^{-1}\|\,\|B^{-1}\|\;\|A-B\|,$ and take $C_{1}\;=\;\|A^{-1}\|\;\|B^{-1}\|.$ $A,B$ are positive definite, denote $\lambda_{\min}(A)\;=\;\text{the smallest eigenvalue of }A,\quad\lambda_{\min}(B)\;=\;\text{the smallest eigenvalue of }B.$ Then $\|A^{-1}\|\;=\;\frac{1}{\lambda_{\min}(A)},\quad\|B^{-1}\|\;=\;\frac{1}{\lambda_{\min}(B)},$ when the norm is the usual spectral/operator norm. Hence, $C_{1}\;=\;\|A^{-1}\|\;\|B^{-1}\|\;\;\leq\;\frac{1}{\lambda_{\min}(A)\,\lambda_{\min}(B)}.$ We have: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$ Substituting this into the expression for $\Delta\mu(\mathbf{x}_{*})$, we obtain:
$|\Delta\mu(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|$
$\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$
@@
$\Delta\sigma^{2}(\mathbf{x}_{*})=\left[k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*\mathcal{D}}\right]-\left[k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*\mathcal{D}}\right].$
$=\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{k}_{*\mathcal{D}}.$
From the mean error analysis, we know: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|,$ for some constant $C_{1}$ that depends on the spectral properties of the matrices. Using the spectral norm (or any subordinate matrix norm), the variance error can be bounded as: $|\Delta\sigma^{2}(\mathbf{x}_{*})|=\left|\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{k}_{*\mathcal{D}}\right|.$ Applying the properties of matrix norms and the Cauchy–Schwarz inequality:
$|\Delta\sigma^{2}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}\|(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|.$
$\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}C_{1}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$

### v2

@@ 6.1 Synthetic Test Problems
Computational Efficiency Analysis. Figure 3 compares the cumulative runtime(in seconds) over $1000$ iterations on low-dimensional Hart6 and high-dimensional Powell50 (results for other test functions are provided in the appendix due to space constraints). In both plots, VecchiaBO incurs a rapidly accelerating runtime, whereas GSSBO and RSSBO remain notably lower than GP-UCB. While VecchiaBO reduces the cost of GP fitting by conditioning on nearest neighbors, its runtime is dominated by the costly maintenance of a structured neighbor graph, which scales poorly with dimensionality and sample size. The running speeds of LR-First, SparseGP, and mini-META are all improved compared to the standard GP-UCB. IPA has no obvious advantage in terms of time consumption because it needs to calculate the sample quality. In contrast, GSSBO and RSSBO cost much less time. Because we use an efficient pilot-based method to compute gradients with minimal computational overhead, the fitting cost per iteration of GSSBO remains around $\mathcal{O}(M^{3})$ once we restrict the active subset to size $M\ll n$. The GSSBO and the random one are often similar in runtime, though the GSSBO can be slightly higher due to the overhead of the gradient-based sample selection. By iteration 1000, the total running time of GSSBO on both functions is only about $10\%$ of that of standard GP-UCB. This advantage of GSSBO becomes more pronounced as $n$ increases.
Figure 2: Cumulative regret of algorithms on synthetic and real-world test problem experiments.
Figure 3: Cumulative time cost of algorithms.
Figure 4: Sensitivity analysis of $Z$.
@@
Optimization Performance Analysis. Figure 2 compares methods on multiple functions, evaluating cumulative regret. Overall, GSSBO achieves comparable performance with the Standard GP-UCB while outperforming the RSSBO. In low-dimensional problems such as Eggholder2 and Hart6, GSSBO has a subtle gap with Standard GP-UCB and VecchiaBO. In contrast, GSSBO significantly outperforms other methods. In high-dimensional settings, such as on Levy20, Powell50, and Rastrigin100, GSSBO achieves the smallest cumulative regret, surpassing other methods. From the experimental results, we can observe that the cumulative regret of our algorithm is sublinear, which is consistent with the theoretical results. In particular, GSSBO achieves these results with a significant reduction in computation time, as shown in Figure 3. GSSBO strikes a combination of performance and efficiency in scalable optimization tasks by maintaining near-baseline regret while significantly improving computational efficiency.
@@
Sensitivity analysis of Hyperparameter $Z$. We further examine how the dynamic buffer parameter $Z$ affects GSSBO. Figure 4 presents results on two functions: Hart6 and Powell50. For RSSBO, $Z$ remains fixed at 4, whereas for GSSBO, we vary $Z\in\{2,4,6,8\}$. On Hart6, larger $Z$ consistently boosts the GSSBO’s performance toward that of Standard GP-UCB, while the RSSBO lags in cumulative regret. Intuitively, for low-dimensional problems, allowing the model to retain more samples helps preserve important information, bridging the gap with the Standard GP-UCB baseline. In contrast, in Powell50, smaller $Z$ leads to slightly better performances for GSSBO, reflecting the benefit of subset updates in high-dimensional landscapes. In general, low-dimensional tasks benefit from a larger $Z$, while high-dimensional problems perform better with more aggressive subset limiting, $Z$ can be effectively tuned to match the complexity of tasks.
@@ 6.2 Real-World Application
To assess the applicability of GSSBO in real-world applications, we applied NAS to a diabetes-detection problem from the UCI repository [47]. We modeled the problem of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ($[32,128],[1e-6,1.0],[1e-6,1.0],[1,8]$ for batch size, Learning Rate, Learning Rate decay, hidden dim), where $y_{t}$ is the test classification error. Each iteration is repeated 5 times. The results in Figure 2 indicate that GSSBO outperforms all competitors.
@@ 7 Conclusion
Appendix A Appendix Theoretical Analysis
@@ A.1 Theorem 1:Analysis on subset GP:
Consider a GP model, where we assume $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime})),$ and given samples $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{N},y_{N})\}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, we have a noise model: $\mathbf{y}=f(\mathbf{X})+\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\sigma_{n}^{2}\mathbf{I}),$ where $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}]^{\top}\in\mathbb{R}^{N\times d}$ and $\mathbf{y}=[y_{1},y_{2},\ldots,y_{N}]^{\top}\in\mathbb{R}^{N}$. The posterior predictive distribution for a test point $\mathbf{x}_{*}$ is Gaussian with the following mean and variance: $\mu(\mathbf{x}_{*})=\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{y},$ $\sigma^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*},$ where $\mathbf{K}$ is the $N\times N$ kernel matrix evaluated at the training samples, i.e., $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$; $\mathbf{k}_{*}\in\mathbb{R}^{N}$ is the vector of covariances between the test point $\mathbf{x}_{*}$ and all training points, i.e., $(\mathbf{k}_{*})_{i}=k(\mathbf{x}_{*},\mathbf{x}_{i})$; $k_{**}=k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the kernel evaluated at the test point itself.
@@
Instead of using all $N$ training samples, consider a subset samples $\mathcal{U}=\{\mathbf{u}_{1},\dots,\mathbf{u}_{M}\}$, where $M\ll N$. Define: $\mathbf{K}_{\mathcal{U}\mathcal{U}}\in\mathbb{R}^{M\times M},\quad\mathbf{K}_{\mathcal{D}\mathcal{U}}\in\mathbb{R}^{N\times M},\quad\mathbf{K}_{\mathcal{U}\mathcal{D}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}^{\top},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ is the kernel matrix among the $M$ inducing points, and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ represents the covariances between the full training samples in $\mathcal{D}$ and the inducing points in $\mathcal{U}$. Using Subset of Regressors (SoR), a low-rank approximation to the kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ is given by: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
@@
$\Delta\mu(\mathbf{x}_{*})=\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{y}.$ (6)
Define the following positive definite matrices: $\mathbf{M_{A}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I},~{}\mathbf{M_{B}}=\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}.$ Then, the difference between $\mathbf{M_{A}}$ and $\mathbf{M_{B}}$ is: $\mathbf{M_{A}}-\mathbf{M_{B}}=\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}.$ We have $\|\mathbf{M_{A}}^{-1}-\mathbf{M_{B}}^{-1}\|\;\leq\;\|\mathbf{M_{A}}^{-1}\|\;\|\mathbf{M_{B}}^{-1}\|\;\|\mathbf{M_{A}}-\mathbf{M_{B}}\|,$ and take $C_{M}\;=\;\|\mathbf{M_{A}}^{-1}\|\;\|\mathbf{M_{B}}^{-1}\|.$ Since $\mathbf{M_{A}},\mathbf{M_{B}}$ are positive definite, denote $\lambda_{\min}(\mathbf{M_{A}})=\text{the smallest eigenvalue of }\mathbf{M_{A}},~{}\lambda_{\min}(\mathbf{M_{B}})=\text{the smallest eigenvalue of }\mathbf{M_{B}}.$ Under the spectral norm, $\|\mathbf{M_{A}}^{-1}\|=\frac{1}{\lambda_{\min}(\mathbf{M_{A}})},\|\mathbf{M_{B}}^{-1}\|=\frac{1}{\lambda_{\min}(\mathbf{M_{B}})}.$ Noting that $\lambda_{\min}(\mathbf{M_{A}})=\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})+\sigma_{n}^{2},~{}\lambda_{\min}(\mathbf{M_{B}})\geq\sigma_{n}^{2},$ we obtain the bound on $C_{M}$ that $C_{M}=\frac{1}{\bigl{(}\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})+\sigma_{n}^{2}\bigr{)}\,\sigma_{n}^{2}}\leq\frac{1}{\lambda_{\min}(K_{\mathcal{D}\mathcal{D}})\,\sigma_{n}^{2}}.$ We have: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{M}\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$ Substituting this into the expression for $\Delta\mu(\mathbf{x}_{*})$, we obtain:
$\begin{split}|\Delta\mu(\mathbf{x}_{*})|&\leq\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;\bigl{\|}(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\bigr{\|}\\[4.30554pt] &\leq\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;C_{M}\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.\end{split}$ (7)
@@
$\begin{split}\Delta\sigma^{2}(\mathbf{x}_{*})&=\bigl{[}\,k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}\bigl{(}\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{k}_{*\mathcal{D}}\bigr{]}-\bigl{[}\,k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}\bigl{(}\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}\bigr{)}^{-1}\mathbf{k}_{*\mathcal{D}}\bigr{]}\\ &=\mathbf{k}_{*\mathcal{D}}^{\top}\bigl{[}\,(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\bigr{]}\mathbf{k}_{*\mathcal{D}}\,.\end{split}$ (8)
From the mean error analysis, we know: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.$ Using the spectral norm, the variance error can be bounded as: $|\Delta\sigma^{2}(\mathbf{x}_{*})|=\left|\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{k}_{*\mathcal{D}}\right|.$ Applying the properties of matrix norms and the Cauchy–Schwarz inequality:
$\begin{split}\bigl{|}\Delta\sigma^{2}(\mathbf{x}_{*})\bigr{|}&\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}\,\Bigl{\|}(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\Bigr{\|}\\ &\leq\|\mathbf{k}_{*\mathcal{D}}\|^{2}\,C_{M}\,\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|.\end{split}$ (9)